OpenStudy (anonymous):

Phillip deposited $5,670 into a savings account 19 years ago. The account has an interest rate of 4.1% and the balance is currently $12,356.04. How often does the interest compound? daily monthly quarterly annually

OpenStudy (anonymous):

quartley probolly

OpenStudy (triciaal):

Accumulated = Principal(1 + rate)^time 12356.04/5670 = (1.041)^t = 2.179 t log 1.041 = log 2.179 t = .338/0.017 = 19.8 in 19 years the interest is compounded 19 times therefore once per year = annually